OpenStudy (anonymous):

Which of the following equations is the slope-intercept form of 4x - 2y = 8? A. y = 2x - 4 B. y = -4x - 8 C. y = 1/2 x + 2 D. y = -2x + 4

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

Which one.

OpenStudy (anonymous):

4x-2y=8 2y = 4x-8 y = 2x-4 which is of the form of y = mx + c where m = slope and c = intercept so its A

OpenStudy (anonymous):

@ashleyD143 right???????????

OpenStudy (anonymous):

yes.